### v2

# What is emergence, after all?

Abstract
The term emergence is increasingly used across scientific disciplines to describe phenomena that arise from interactions among a system’s components but cannot be readily inferred by examining those components in isolation. While often invoked to explain higher-level behaviors—such as flocking, synchronization, or collective intelligence—the term is frequently used without precision, sometimes giving rise to ambiguity or even mystique. In this perspective paper, we clarify the scientific meaning of emergence as a measurable, physically grounded phenomenon. Through concrete examples—such as temperature, magnetism, and herd immunity in social networks—we review how collective behavior can arise from local interactions that are constrained by global boundaries. By disentangling emergence from vague overuse, we emphasize its role as a rigorous tool for understanding complex systems. Our goal is to show that emergence, when properly framed, offers not mysticism but insight.
keywords:
Emergence $|$ Reductionism $|$ Complex Systems $|$ More is Different
[…]
This manuscript was compiled on August 11, 2025
[…]
Emergence occurs where a whole exhibits properties absent from its individual parts—an idea that traces back to Aristotle’s dictum that “the whole is something besides the parts” (1). In science, the term refers to patterns, structures, or behaviors that are not straightforwardly deducible from components in isolation (2, 3, 4, 5, 6, 7), yet are often captured by coarse, predictive descriptions that do not track every micro-detail. You can see emergence in the wild—flocks wheeling as one, fireflies locking into rhythm, insect societies coordinating without commanders (8)—and on through computer simulations, where minimal rules produce striking order: Reynolds’ “boids” (9), cellular automata (10), and the Game-of-Life glider (11).
Philosophical roots of the concept trace back to Mill’s distinction between additive “mechanical” effects and qualitatively new “chemical” ones (12). Lewes later coined the term “emergent” for such cases (13), and early British emergentists (14, 15, 16) developed it into a layered view of nature in which higher-level laws complement, and sometimes influence, lower-level processes. This tradition faded as reductionism gained ground, driven by the intertheoretic reduction models (17, 18, 19). The debate revived with arguments for multiple realizability (20, 21, 22) and Anderson’s claim that “more is different” (23), followed by work on supervenience and causal autonomy (24). In recent years, the term has evolved from describing specific transitions, such as the “emergence of metabolism” (25), to becoming a popular buzzword, gaining more usage in scientific and non-scientific discourses (26, 23). Fig. 1 shows this trend.
Here we re-anchor emergence in measurable, physically grounded mechanisms. We begin with textbook cases from physics, where coarse-graining and symmetry breaking compress microscopic dynamics into macroscopic laws. We then turn to biology and the social sciences, where collective behavior arises from local interactions. Finally, we consider emergence in networks, showing how the same conceptual toolkit applies across domains. We aim to show that emergence is about an economy of description that discards inessential detail while preserving predictive power.
Figure 1: Usage trajectories of key concepts in complexity discourse, 1900 – 2022. Per-million-word frequency of six terms in the Google Books English corpus, 1900–2022. Seven-year–smoothed curves (smoothing = 3) reveal the rise of “emergence” and “emergent” compared to other famous keywords based on data downloaded via Google Ngram Viewer (27).
Level of Description
Knowledge unfolds across successive levels of abstraction—from raw experience to language and higher-order concepts (28). Confusion between these levels distorts understanding. Emergence is everywhere, and the surprise it evokes often reflects where we choose to stop asking more questions (29, 24). What we call “emergent” typically depends on the limits of our knowledge, tools, or perspective—in other words, on epistemology (30, 31). Mixing red and green paint yields a dull brown—basic chemistry explains it well enough for most, and few would call the result emergent. Yet a physicist who tracks the same blend down to quantum-mechanical interactions would hardly find the task trivial.
Every phenomenon can be analyzed at multiple descriptive levels, and what feels emergent at one level may be obvious at another (30). Explanation is always a matter of perspective: an economist analyzing a financial crisis need not invoke atoms or molecules but instead works at the social scale relevant to the questions at hand. Even so, emergent behaviors are not merely in the eye of the beholder—they can display objective, quantifiable signatures such as information flow or causal strength (32, 33). We can make emergence more rigorous by formalizing it through coarse-graining maps that discard detail yet preserve prediction.
Figure 2: Coarse-graining with different maps.
(a) A global coarse-graining that yields temperature $T$ by averaging the kinetic energy of all particles, assuming local equilibrium; $\mathcal{F}:\bigl{(}v_{1},\cdots,v_{N}\bigr{)}\;\longmapsto\;T.$ (b-c) Local majority-vote map applied to blocks of four spins $(s_{i},s_{j},s_{k},s_{l})$, returning a block spin $S_{I}$ defined as: $S_{I}=\operatorname{sign}\left(s_{i}+s_{j}+s_{k}+s_{l}-2\right)$, where $\operatorname{sign}(x)$ is 1 for $x>0$, and -1 otherwise; $\mathcal{F}:\bigl{\{}(s_{i},s_{j},s_{k},s_{l}\}\bigr{)}\;\longmapsto\;\{S_{I}\}.$ (c-d) The same map is applied to the previously coarse-grained block. Applying $\mathcal{F}$ on block (d) will result in a single upward block spin.
Many-to-one Maps & Coarse-graining
Emergence is present when there exists a many-to-one map from a micro-level theory (more fundamental, detailed, or lower-level) to a macro-level theory (higher-level), such that the macro description remains predictive even after discarding most of the microscopic detail (34). As Fig. 2 presents, a map is a selective, structured, purpose-driven abstraction—useful precisely because it is not the whole territory it depicts (28, 35). Every map omits, distorts, or abstracts; acknowledging those sacrifices is part of scientific hygiene (35). Coarse-graining leaves us with just enough structure and real patterns to construct an autonomous and useful theory (36). This is why engineers who build bridges do not need to take a course in quantum field theory, or why airplanes can fly safely despite our incomplete understanding of quantum gravity.
Coarse-graining is the process of identifying which features of a system are essential for capturing its macroscopic behavior and which can be systematically discarded. Let the system’s state at time $t$ be given by the vector $\mathbf{x}(t)=\bigl{(}x_{1}(t),\dots,x_{N}(t)\bigr{)}$, where a high-resolution dataset comprises $N$ time series, and each $x_{i}(t)$ represents the evolution of the $i$-th microscopic variable. We can then formally view coarse-graining as a process of lossy data compression, which can be described as a mapping:
$\mathcal{F}:\;\bigl{(}x_{1}(t_{i}),\dots,x_{N}(t_{k})\bigr{)}\;\longmapsto\;\bigl{(}X_{1}(t^{\prime}_{j}),\dots,X_{N^{\prime}}(t^{\prime}_{\ell})\bigr{)},$ (1)
where $N^{\prime}<N$, and the coarse-grained observables $\{X_{j}\}$ may be defined on a different (typically coarser) time grid $\{t^{\prime}_{\ell}\}$. Fig. 2 shows two different sets of coarse-graining where they simply average out the observables. The key practical questions here are whether there exists a finite set of macroscopic parameters that can adequately describe the phenomena of interest within a specified tolerance, and if so, which mapping $\mathcal{F}^{\star}$ minimizes the chosen distortion measure while achieving that macro-level description. The outcome is a systematic and metric-aware approach to transition from micro-level dynamics to macro-level dynamics.
The very notion of a map between levels of description tempts us to believe that deeper, more fundamental layers offer better explanations—an intuition reinforced by Physics’ success in reducing systems to atoms and beyond. But how far should we go? What counts as the most fundamental level?
Reductionism & “Theory of Everything”
Physics is expressed in languages that are intrinsically redundant—coordinates, gauge potentials, duality frames, and so on—and those redundancies shift as we move between scales (37). What we call “fundamental”—even spacetime and gauge fields—may not exist in a more microscopic theory, but rather emerge after coarse-graining. Spacetime geometry and general relativity might also be emergent, arising from quantum theory through entanglement and boundary dynamics (37). This realization, however, does not make emergence a “secret sauce.” It simply shows how nature organizes complexity through scale, and “fundamental” is often a provisional concept.
Each scale brings qualitatively new behaviors that demand their own inquiry. So, reducing a phenomenon to fundamental laws does not mean we can reconstruct everything from them. Reductionism is not wrong—it is just insufficient for understanding the universe. Even a complete microscopic “Theory of Everything” would leave many of the essential rules governing higher-level systems untouched (38). Macroscopic behavior is determined by emergent parameters that are largely insensitive to the fine details of the underlying microscopic governing equation (39). Unifying quantum mechanics and gravity would undoubtedly be a milestone in Fundamental Physics, but it would not immediately explain why financial markets crash. These are collective phenomena with different ontologies, which cannot be derived by reduction alone (38).
A New Ontology
In Classical Mechanics, the complete state of a system is described by the positions and momenta of its particles (40). Concepts like temperature or pressure only become meaningful in the thermodynamic limit—for instance, when considering systems with $10^{23}$ particles (41). Large language models (LLMs) display a digital analogue of this behavior: they acquire new capabilities, such as multi-digit arithmetic or spatial reasoning, only after reaching a sufficient scale (42). In this way, temperature emerges as a qualitatively different property in the large-size limit. Therefore, Thermodynamics and Classical Mechanics do not share the same conceptual structure and notion or ontology (43). It makes no sense to assign a temperature $T$ to a single particle since it is a macroscopic property of a bulk system, not of particles. Temperature is a local and direct emergent property (34): local, because the temperature at any point depends only on particles within a nearby volume—not on distant parts of the system—thus preserving the spatial locality of the underlying theory; and direct, because the coarse-graining map is a simple analytic function—essentially the mean kinetic energy per particle, $T\propto\langle v^{2}\rangle$—rather than an algorithmically complex lookup.
Classical Mechanics itself emerges from a more fundamental quantum theory, which is, in turn, ontologically different. In Classical Mechanics, a particle is a point in phase space with well-defined position and momentum. In Quantum Mechanics, the ontology shifts: a particle is represented by a vector in Hilbert space, and is described by a wavefunction (44). Position and momentum are no longer coordinates but observables. The uncertainty principle ensures that no wavefunction can define both position and momentum with perfect precision. Thus, quantum mechanics offers no definitive answer to the classical question of “where” a particle is, or “how fast” it is moving (44). Nevertheless, the classic macroscopic theory remains applicable in its domain, and it is consistent with the underlying quantum description through the processes of decoherence and measurement, which effectively produce the appearance of wavefunction collapse (45).
Effective Theories
Since the advent of Statistical Mechanics, we have learned how coarse-graining microscopic degrees of freedom gives rise to macroscopic quantities like temperature or pressure, and how these quantities are connected through equations of state (46). Statistical mechanics not only explains why thermodynamics works but also defines its domain of applicability. For example, the domain of applicability of temperature can be expressed as
$T\;=\;\frac{1}{3}m\,\langle v^{2}\rangle\!\left(1\pm\frac{1}{\sqrt{N}}\right),$ (2)
where $m$ is the particle mass, $\langle v^{2}\rangle$ is the ensemble-averaged squared speed, $N$ is the number of particles, and the Boltzmann constant is set to one, $k_{B}=1$ (47). For large $N$, the fluctuation term $1/\sqrt{N}$ vanishes, making temperature a sharply defined coarse-grained variable. For small systems, however, fluctuations are significant, and temperature loses descriptive power. Therefore, in the large-size limit, Thermodynamics is not merely an approximation to statistical mechanics—it is an effective theory (48), one that captures the essential behavior of macroscopic systems using a few key variables at the right scale (49, 50). Instead of tracking the detailed positions and velocities of every particle, we use variables like temperature, pressure, and entropy, governed by their own simple and powerful laws at the level of macrostates (51, 52). Effective theories isolate what matters at a given scale and discard what does not (53). They are not shortcuts; they are self-contained, predictive, and often universal descriptions of emergent levels of reality. A mean-field theory is the zeroth-order effective description (54). It is invaluable as a first cut, exact when fluctuations vanish, but insufficient whenever correlations hold the key to collective behavior (55).
Yet not all effective theories are created equal. In some systems—particularly those involving biological, cognitive, or social dynamics—the coarse-grained variables do not merely summarize the current state of the underlying components. They also retain memory, storing information about past configurations. This historical dependence alters the structure of the effective theory, introducing higher-order terms that reflect feedback, path dependence, and self-reference (34). Such systems are still governed by physical laws, but their dynamics can become computationally intractable, meaning that no tractable microscopic derivation will fully recover the macroscopic behavior (56). The emergence of life, mind, or social structure appears puzzling not because they break with physicalism, but because they represent a qualitatively more complex class of emergent phenomena (56). The first step toward understanding how emergence occurs is to identify when and under what conditions it arises, as well as its immediate consequences.
Onset of Emergence & Symmetry Breaking
A clear and intuitive example of how emergence happens is magnetization. In a fridge magnet, the collective alignment of billions of electron spins produces a macroscopic magnetic property, denoted $m(T)$, which appears only when the system is below a critical temperature $T_{c}$. For a fridge magnet, it is around $450^{\circ}$C. Heating the material beyond this point destroys the alignment, and the magnetic property disappears in a continuous phase transition (47). A single governing equation (Hamiltonian) describes both magnetized and non-magnetized states, and at the critical point $T=T_{c}$, the spin-flip symmetry of the Hamiltonian remains intact, but the actual state chooses a specific direction, resulting in $m\neq 0$ (54). This symmetry breaking defines an order parameter that compresses the full $10^{23}$-spin microstate into a single coarse-grained vector whose dynamics—like domain walls or spin waves—obey new effective laws (57, 58, 23).
At the critical point, key properties of the system—such as correlation length and relaxation time—diverge (59). These divergences, characterized by critical exponents, serve as clear markers of the onset of emergence. More interestingly, different systems composed of distinct elements can exhibit the same critical behavior (60). That is, systems with very different microstructures can fall into the same universality class and be described by the same macroscopic theory. This substrate-independence also means that from the perspective of the emergent behavior, one cannot tell which microscopic system produced it. The liquid–gas transition belongs to the same universality class as the ferromagnetic transition. Universality classes, therefore, organize not just critical exponents but also the very symmetry content that survives at macroscopic scales, reinforcing why broken symmetry is the natural language of emergence (23, 37). In this sense, “more is different” (23): once many degrees of freedom lock together, they generate effective laws—and sometimes entirely new effective symmetries—that are absent from, yet fully compatible with, the underlying dynamics.
Broken symmetries underpin certain emergent behaviors (23, 61, 62, 63), but not all. For example, it is not straightforward to extend the statistical mechanics framework of critical phenomena to describe abrupt qualitative changes in dynamical systems. Likewise, the emergence of life from chemical interactions does not neatly correspond to a phase transition in the physical sense, despite claims to the contrary by some physicists (64). Intricate phenomena such as consciousness or thought could, in principle, arise from entirely different physical substrates—biological neurons or silicon circuits alike (65, 66)—provided there is sufficient evidence that they result from well-defined critical phenomena (67).
Universality & Dualities
Technically, critical universality manifests only in the critical region around a continuous phase transition: as the system approaches the transition, the renormalization-group flow is drawn toward a scale-invariant fixed point that is attractive along irrelevant directions but unstable along relevant ones, making macroscopic behavior largely independent of microscopic details (61). However, this is not the only notion of equivalence and universality in Physics.
Dualities show that theories with ontologically distinct variables—bosons versus fermions, or gravity versus no gravity—can represent the same underlying quantum physics. In $1+1$ dimensions, the bosonic sine–Gordon model is exactly dual to the fermionic Thirring model, exchanging solitonic waves for interacting particles (68). In AdS/CFT, a conformal field theory without gravity on a $d$-dimensional boundary is holographically equivalent to a $(d+1)$-dimensional bulk theory with gravity (69). Even composite objects such as hydrogen atoms—each built from two spin-$\frac{1}{2}$ fermions—can act as effective bosons. Therefore, the large-scale description is not dictated by microscopic “stuff,” but by deeper symmetries and consistency conditions, reminding us that macro structures can arise as alternative micro read-outs of the same physical substrate.
Figure 3: Emergence of herd immunity in social networks. (a–c) Each immune node (blue) deactivates its adjacent edges, forming a local “firewall” that halts transmission to nearby susceptible nodes. The more connections an immune node has, the larger the firewall it establishes. As immunization progresses, these firewalls begin to overlap and coalesce, forming a larger collective barrier that expands nonlinearly—often outpacing the fraction of immunized individuals. (d) Once a critical fraction of nodes (white) is immunized, herd immunity emerges: the remaining susceptible nodes (black) become indirectly protected. This emergent protection is shaped by both the structural and geometric properties of the social network. (e) The thick solid curve shows the remaining susceptible fraction $\pi_{\mathcal{S}}$ as a function of the immunized fraction $\pi_{\mathcal{R}}$, constrained by $\pi_{\mathcal{S}}+\pi_{\mathcal{R}}\leq 1$. The shaded region quantifies the share of individuals shielded through structural (indirect) immunity. Horizontal intersections of the curve indicate total immunity thresholds, with $\pi^{*}_{\mathcal{R}}$ marking the structural herd immunity point. (f) The network in panel (d) can be rearranged to reveal an interface of $\mathcal{SR}$ links separating immune (white) from susceptible (black) nodes. The density of these interface links, $\rho_{\mathcal{SR}}$, serves as a proxy for the potential of epidemic containment and the strength of indirect protection.
Emergence in Networks
To move beyond examples from physics, we now examine how large-scale patterns can emerge from local interactions in social and informational systems. Consider a set of $N$ points where each pair is connected with probability $p$. This defines an Erdős–Rényi random graph with $N$ nodes and roughly $p\binom{N}{2}$ links or connections. As $p$ increases, the network shifts from isolated nodes ($p=0$) to a fully connected graph ($p=1$), with clusters forming in between. Given the value of $p$, it is interesting to look at the behavior of the cluster with the most number of connected nodes, the giant component. Letting $\sigma$ be the fraction of nodes it contains, we find that as in the large-size limit ($N\to\infty$), the network undergoes a phase transition such that the size of its giant component is zero below a critical value $p_{c}$, and positive above it. Here, $p$ is the control parameter, and $\sigma$ represents the order parameter, mirroring magnetization in ferromagnets, which appears below the critical temperature. When $p>p_{c}$, large-scale connectivity emerges, and the network is said to percolate (70).
The emergence of the giant component is foundational: the Internet stays functional despite random failures because its giant component remains intact (70). Likewise, in epidemic dynamics, if an infected person enters a social network and contacts someone in the giant component, the disease can potentially spread to nearly everyone after enough time. In this way, the onset of an epidemic reflects the same percolation principles that govern network connectivity (71, 72). We say that an outbreak occurs or an epidemic emerges when a significant number of individuals in a population become infected. The onset, size, and time scales of this emergence can be measured and modeled through the contagiousness of the disease and how the contact network is structured and changes over time (73). The emergence of large-scale connectivity is one example among many in networks. Furthermore, it is also easy to argue that communities emerge when networks evolve with local rules (74).
Emergence of Herd Immunity
Vaccination is a key strategy for suppressing epidemics. Take measles, a highly contagious disease with a basic reproduction number of $R_{0}\sim 15$—meaning one case can infect around fifteen others early on in a fully susceptible population (75, 76). Before routine vaccination in the UK began in the late 1960s, outbreaks recurred every 2–3 years as the number of susceptible children quietly rose above the transmission threshold.
Vaccination protects in two ways: directly, by immunizing individuals and removing them as nodes from the transmission network, and indirectly, as Fig. 3(a-c) shows, by forming a firewall (or firebreak (77)) and breaking the chains along which the infection might spread (78). In a vaccinated population, before the introduction of a new infection, individuals are either susceptible ($\mathcal{S}$) or immune ($\mathcal{R}$). The network representation of this population can always be rearranged to separate these two groups, with the interface formed by links connecting $\mathcal{S}$ and $\mathcal{R}$ nodes (Fig. 3f). The size of the immune set represents the direct benefit of vaccination, while the number of interface links serves as a proxy for indirect protection.
When enough immune individuals form clusters, they fragment the transmission network into disconnected parts (Fig. 3d), stopping the infection from spreading widely. In this way, immunity percolates through the population: local outbreaks fail to propagate beyond their immediate neighborhood, and as a result, even unvaccinated individuals benefit from this collective protection, known as herd immunity (70). The extent of the emergent barrier—the firefront—is quantified by the density of susceptible-immune ($\mathcal{SR}$) links, $\rho_{\mathcal{SR}}$. Each $\mathcal{SR}$ link both blocks transmission and reduces the downstream branching factor of the pathogen, making $\rho_{\mathcal{SR}}$ a nonlinear, structural measure of resistance. The question, then, is not just how many people must be vaccinated, but who should be vaccinated with what order for herd immunity to emerge?
In a hypothetical fully mixed population—where each individual interacts with all others with equal probability—this question has a simple answer. If a fraction $\pi_{\mathcal{R}}$ of the population is randomly vaccinated before the introduction of the disease, then a typical infected person can effectively infect $R_{e}=(1-\pi_{\mathcal{R}})R_{0}$ more people. Outbreaks become unsustainable when $R_{e}<1$, yielding the classical threshold. $\pi_{\mathcal{R}}^{*}=1-\frac{1}{R_{0}}.$ This formulation, however, assumes that immunity acts uniformly and independently across individuals, neglecting the underlying contact structure of real populations (78). However, real populations form heterogeneous, spatially embedded networks (79). Most interactions cluster within social, geographic, or institutional contexts (80).
A node’s number of connections is referred to as its degree (70). Immunizing a highly connected node can disrupt many transmission routes, whereas vaccinating randomly chosen low-degree nodes typically offers little indirect protection. Yet, the combined impact of removing multiple nodes is not easy to predict: the strength of the resulting indirect immunity depends sensitively on the network’s spatial structure and the distribution of immunized nodes within it (81). As Fig. 3(a-c) shows, each immune node acts as a local firewall, suppressing links through which infection would otherwise spread. As immunity accumulates, these firewalls overlap, and their collective effect can grow faster than the immunized fraction (Fig. 3d). By immunizing a fraction $\pi_{\mathcal{R}}$ of the population, at most a fraction $\pi_{\mathcal{S}}\leq 1-\pi_{\mathcal{R}}$ remains susceptible. The exact value of $\pi_{\mathcal{S}}$ depends on the network’s structure and how immunity fragments it. The difference, $1-\pi_{\mathcal{R}}-\pi_{\mathcal{S}}$, corresponds to individuals who are not vaccinated but are indirectly protected, represented by the hatched area in Fig. 3e.
Recent studies formalize this mechanism using bond percolation and message-passing on real and synthetic networks (81, 79). They track the dynamics of the susceptible giant component and the firefront density, $\rho_{\mathcal{SR}}$, across diverse mixing patterns. These analyses reveal a competing mechanism: targeting superspreaders or implementing acquaintance immunization increases collective immunity, while the spatial structure of the network can hinder it by localizing the protection, leaving remote, susceptible pockets vulnerable (81).
Herd immunity represents a paradigmatic case of emergence in social networks; Local immunizations combine nonlinearly to produce a global shield shaped by the network’s structure. Recognizing this not only deepens our understanding of collective behavior but also equips us with practical tools to design smarter, more efficient epidemic interventions.
Weak vs. Strong Emergence
Finally, the consistency between theories with different ontologies raises the question of how much causal autonomy higher levels truly possess—a distinction philosophers frame as weak versus strong emergence. Scientific communities use the term “emergence” to indicate that some collective behaviors have explanatory autonomy (21) and are, in principle, possible but practically difficult to derive from purely microscopic descriptions, even when all micro-level details are known (4). Philosophers often refer to this as weak emergence (82), while leaving room for the idea of strong emergence—where some phenomena might be fundamentally irreducible and governed by distinct principles (83). Strong emergence suggests that nature might include layers that follow their own fundamental rules—possibly violating physical laws, creating inconsistencies between micro and macro theories, or breaking causal closure or locality. From the standpoint of science, there is no mystery or divinity in emergence. If a property is measurable, it is physical, and any scientific explanation must reference the constituents and their interactions, no matter how surprising the outcome appears.
In scientific practice, “emergence” almost always refers to weak emergence: a phenomenon grounded in the system’s components and interactions at the right scale, with no need for non-physical explanations. One can argue that a microscopic theory is incomplete or inapplicable to some macroscopic regimes—but to invoke strong emergence in domains like consciousness or social behavior (84, 85) amounts to introducing new causal principles not anchored in substrate dynamics, and thus steps outside the scope of scientific method (86). Therefore, if a macro-level description effectively explains a phenomenon—even more parsimoniously, with fewer micro-level details than a microscopic one—nothing beyond physicalism is at play. There have been, of course, historical movements that made strong-sounding claims about emergent phenomena without yielding testable or predictive frameworks (87, 88, 89, 90, 91, 92).
As causality connects events within the same scale, emergence connects descriptions across scales, revealing how collective patterns arise from local interactions. But the influence does not only run upward. Much like the walls of a container shape the motion of the particles inside, higher-level structures can constrain or guide the behavior of their components without violating the underlying laws. This interplay is often described as downward or top-down causation (93, 94).
It is tempting to wonder whether complex aesthetic outcomes might also emerge from simple, knowable constraints—a question deeply resonant with the concept of emergence itself. If Poe’s self-portrait in The Philosophy of Composition (1846) is taken at face value, at least one kind of aesthetic complexity is not strongly emergent; it can be reconstructed from an explicit, step-wise recipe. Poe famously crafted the poem The Raven by first fixing a desired emotional effect—melancholy—and then deliberately selecting every element (theme, meter, and the refrain “Nevermore”) to achieve that goal (95). He insists good writing is methodical and analytical, explicitly rejecting artistic intuition. While many, including T. S. Eliot, argue that something ineffable lies beyond algorithmic methods (96), Poe’s essay reminds us that not all creativity is opaque to analysis—even if many artists and readers prefer it that way. Today, sentiment-analysis tools can map—and perhaps replicate—a narrative’s emotional arc (97, 98), sharpening our insight into how writers engineer intricate effects. However, until we can define such processes rigorously and run controlled experiments on well-defined creative variables, any final judgment about whether art is ultimately rule-bound or irreducibly spontaneous must remain tentative.
Conclusion and Discussion
There is a story in Persian about a modest art instructor who could skillfully draw many animals—rabbits, deer, birds—but always avoided one: the horse. One day, the students insisted that he draw a horse. Reluctantly, he began from the head, moved gracefully down the body, but as he reached the legs and hooves—his weak point—he hesitated. Then, with a swift stroke, he drew tall grass over the lower legs, hiding the part he could not render. When the students asked him why he added grass, he simply replied that horses naturally belong in fields, neatly sidestepping the truth. In many scientific papers, the term emergence is treated as it was with the hooves. The arguments are precise and formal until they reach a point that resists proper modeling or measurement, at which point people resort to vague arguments to cover up for it. Lacking a consensus on how emergence should be defined or observed as a physical quantity, authors often cloak conceptual gaps with terms that sound profound but remain undefined. Such ambiguity risks obscuring what is truly being explained.
Science is compatible with a form of pluralism that affirms the reality of higher-level causal powers (99). As scientists, we must use language carefully and remain grounded in physical mechanisms. The word emergence is highly used in the field of complex systems. While vague or mystical references to emergence may sound compelling, selling complexity science by mystifying emergence and invoking some spiritual dimensions is a great disservice (100). We seek to clarify emergence by framing it as a many-to-one coarse-graining process that retains predictive power while discarding most micro-level details. Within this framework, we differentiate between weak emergence, which is fully compatible with underlying dynamics, and the more philosophically problematic notion of strong emergence. Furthermore, we mention how measurable quantities—such as information flow, causal strength, order-parameter dynamics, and symmetry breaking—can transform emergence from mere rhetoric into testable science. Finally, as a concrete example in a more social setting, we mentioned recent work on herd immunity in structured contact networks where emergence can be rigorously defined, quantified, and linked to network geometry. Understanding the emergence of herd immunity highlights that epidemic control, like many collective phenomena, cannot be fully understood without considering the structure and correlation across social scales.
[…]
AKR is grateful to Petter Holme and Yasser Roudi for their thoughtful discussions and acknowledges support from the Independent Research Fund Denmark (EliteForsk grant to Sune Lehmann), the Carlsberg Foundation (The Hope Project), and the Villum Foundation (NationScale Social networks). For this manuscript, grammar and style improvements were aided by AI language tools, under the author’s full supervision and approval.
[…]